OpenStudy (anonymous):

A car is traveling at 50.3 mi/h on a horizontal highway. The acceleration of gravity is 9.8 m/s 2 . If the coefficient of friction between road and tires on a rainy day is 0.064, what is the minimum distance in which the car will stop? (1 mi = 1.609) Answer in units of m

OpenStudy (kittiwitti1):

We just learned this in class so I'm not familiar with this kind of problem but... this is as close a question as I could find. Since it's not exactly the same I can't help you solve it all the way, sorry.\[\Sigma F _{y} = N - F _{g} = ma\][ ma ] cancels out to 0\[N = F _{g}\]\[\Sigma F _{x} = F _{A} - f = ma\][ ma ] cancels out to 0\[F _{A} = f = \mu N\]\[\mu x = \frac{ F _{A} }{ N } = \frac{ F _{A} }{ F _{g} }\] After that use [mu]x to find Fa (mu is that weird Greek "m" letter in case you didn't know):\[F _{A} = \mu xN\] I'm guessing you can use that to find acceleration and in turn use kinematics to figure out the answer...

OpenStudy (kittiwitti1):

Tagged my friends here - @AllTehMaffs @Kfins99 @calculusxy @nincompoop @Zale101